# MAGIC-VQA: Multimodal And Grounded Inference with Commonsense Knowledge for Visual Question Answering

### v1

@@ Abstract
Visual Question Answering (VQA) requires reasoning across visual and textual modalities, yet Large Vision-Language Models (LVLMs) often lack integrated commonsense knowledge, limiting their robustness in real-world scenarios. To address this, we introduce MAGIC-VQA, a novel framework that enhances VQA by systematically integrating commonsense knowledge with LVLMs. MAGIC-VQA employs a three-stage process: (1) Explicit Knowledge Integration from external sources, (2) By-Type Post-Processing for contextual refinement, and (3) Implicit Knowledge Augmentation using a Graph Neural Network (GNN) for structured reasoning. While GNNs bring greater depth to structured inference, they enable superior relational inference beyond LVLMs. MAGIC-VQA bridges a key gap by unifying commonsensse knowledge with LVLM-driven reasoning, eliminating the need for extensive pre-training or complex prompt tuning. Our framework achieves state-of-the-art performance on benchmark datasets, significantly improving commonsense reasoning in VQA.
@@
Shuo Yang1, Siwen Luo2, Soyeon Caren Han1, Eduard Hovy1 1The University of Melbourne,2The University of Western Australia
@@ 1 Introduction
Visual Question Answering (VQA) Antol et al. (2015); Goyal et al. (2017); Yue et al. (2024) is a complex task requiring models to understand the interaction between visual inputs and textual queries. In recent years, Large Vision-Language Models (LVLMs) Dai et al. (2023); Xue et al. (2024); Wang et al. (2024b); Chen et al. (2023, 2024b); Liu et al. (2024b); Li et al. (2024); OpenAI (2024b); Gemini Team (2024) have made substantial progress in VQA through extensive pre-training on massive image-text datasets and instruction tuning. These models excel at object-level visual recognition and semantic understanding, capturing attributes such as spatial relationships and contextual details.
Nevertheless, LVLMs often face challenges on questions requiring commonsense reasoning—particularly those hinging on implicit contextual cues or everyday world knowledge Zhou et al. (2023); Ye et al. (2022); Li et al. (2023b) [^1] The sample illustrations can be found in Section 6 and Appendix.. To overcome this limitation and improve the performance on commmonsense VQA, different methods have been explored. For example, Multimodal retrieval-augmented generation leverages dense retrieval to inject external multimodal information into the generation process, thereby enhancing the factual grounding of LVLMs Lin and Byrne (2022); Hu et al. (2023). Multimodal prompt tuning harnesses the model’s innate commonsense knowledge by carefully crafting prompts that combine visual and textual cues from representative samples, guiding LVLMs to leverage their internal reasoning for context-rich answers Wei et al. (2022); Zhang et al. (2023). However, static prompt design usually lacks the dynamic adaptability required for novel scenarios, resulting in limited generalization to unseen or diverse inputs. Additionally, graph-based approaches utilize Graph Neural Networks (GNNs) to incorporate structured commonsense knowledge Ravi et al. (2023); Wang et al. (2022), which surpasses the limitations of purely parametric LVLMs, enabling models to capture explicit and implicit knowledge connections via structured graphs.
@@ 2.1 VLPM and LVLMs on VQA
Vision-Language Pretrained Models (VLPMs) like ViLBERT Su et al. (2019), ALBEF Li et al. (2021) and VILT Kim et al. (2021) have advanced Visual Question Answering (VQA) by improving the alignment between visual and textual modalities in the last few yearsLong et al. (2022a). Recently, Large Vision-Language Models (LVLMs) like InstructBLIP Dai et al. (2023), LLaVA Liu et al. (2024b),GPT4o OpenAI (2024b) and Gemini1.5 Gemini Team (2024) further push the boundary of VQA with strong in-context learning capability through extensive pre-training and instruction-tuning on large-scale image-text datasets. However, these models still face challenges with questions requiring commonsense knowledge that is intuitive and straightforward for humans, such as reasoning based on implicit contextual cues or general world knowledge Ye et al. (2023); Wang et al. (2023); Chen et al. (2024a); Yang et al. (2025). The resource-intensive nature of these models further makes it infeasible to train a model from scratch specifically for enhanced commonsense understanding.
@@ 2.2 Commonsense Knowledge Integration for Visual Question Answering
Several studies have highlighted the critical role of commonsense knowledge integration in enhancing the performance of VLPMs and LVLMs on VQA tasks Wu et al. (2022); Zhang et al. (2022); Ding et al. (2022); Wang et al. (2024c). These methods can be classified into two approaches: explicit commonsense knowledge integration and implicit commonsense knowledge integration.
1) Explicit commonsense knowledge integration directly incorporates external commonsense knowledge into model training through instruction tuning or prompt tuning. For example, VLC-BERT Ravi et al. (2023) encodes the contextualized commonsense knowledge of the question phrases as additional textual features and integrates with object visual features to fine-tune the VL-BERT Su et al. (2019). MM-CoT Zhang et al. (2023), T-SciQ Wang et al. (2024a) and KAM-CoT Mondal et al. (2024) fine-tune models on commonsense-augmented Chain-of-Thought (CoT) data to enhance their reasoning processes. However, these methods suffer from static commonsense integration without dynamic filtering to adjust to varying input contexts, resulting in potential noise that impedes nuanced reasoning.
2) Implicit Commonsense Knowledge Integration focuses on distilling knowledge from a teacher to a student model without directly incorporating external datasets. For example, Dai et al. (2022) distill knowledge from the dual-stream VLP model CLIP Radford et al. (2021) into BART Lewis (2019), achieving strong zero-shot performance on VQA. Park et al. (2024) proposed a novel method to distill knowledge from LLMs focusing on specific image regions, then guiding the LLM to infer commonsense knowledge about those areas. These methods often overlook the structured interplay among visual, textual, and commonsense cues, limiting their ability to perform nuanced reasoning.
@@ 3.4 Commonsense Grounded Inference
In the final inference stage, we combine all processed elements—original inputs ($I,Q,C$), refined commonsense triplets (with assigned relevance levels) and GNN-generated confidence scores—into a unified input structure for inference with LVLMs [^3] A complete input example is provided in Appendix G. By fusing explicit and implicit commonsense knowledge with visual and textual signals, the LVLMs can reason effectively about nuanced relationships, delivering answers better aligned with real-world understanding.
@@ 4.1 Dataset
Figure 3: The distribution of categories of commonsense knowledge. CS-PE refers to physical entity-related commonsense, CS-EC to event-centered related commonsense, and CS-SI to social interaction-related commonsense.
@@ 4.2 Commonsense Knowledge Distribution
To tailor the commonsense knowledge to each dataset’s specific reasoning requirements, we analyze the distribution of commonsense types across each dataset using GPT4 OpenAI (2024a) [^4] The prompt template is in Appendix H. As Figure 3 suggests, ScienceQA requires more Physical Entity (CS-PE) knowledge, possibly due to its focus on factual and procedural scientific concepts. Meanwhile, TextVQA, which often involves contextual understanding in images, benefits more from Event-Centered (CS-EC) knowledge. MMMU, however, requires a balanced mix of Physical Entity, Event-Centered, and Social Interaction (CS-SI) commonsense due to its multi-disciplinary nature. As a result, we set the by-type filtering ratio of {CS-PE:CS-EC:CS-SI} mentioned in Section 3.2 as {0.7:0.15:0.15} in ScienceQA, {0.2:0.6:0.2} in TextVQA,{0.33:0.33:0.33} in MMMU dataset.
@@ 5.1 Explicit Commonsense Knowledge
We evaluated the explicit integration of commonsense knowledge triplets by systematically testing four configurations of: (1) None: Inputs with no additional commonsense; (2) CS Sources: Inputs augmented with commonsense from questions, images, or captions; (3) CS Categories: Inputs augmented with commonsense grouped by category (Physical Entities, Event-Centered, Social Interaction); and (4) All CS: Inputs augmented with all retrieved commonsense  [^5] Each experiment is tested with a fixed number of $k=6$ to maintain a fair comparison.. As in Table 2, integrating explicit commonsense consistently improves performance across all baselines and three datasets. For instance, on ScienceQA, GPT-4O’s accuracy rises from 76.45% (None) to 81.22% (All CS), and Qwen2VL improves from 51.10% to 57.42% on MMMU under the same setup. Examining the effect of source-based commonsense reveals that image-driven knowledge (CS-I) typically provides the largest gains. For example, LLaVA1.6 on MMMU jumps from 48.38% to 53.52% with CS-I, surpassing the minor improvements from CS-Q or CS-C. This suggests that leveraging image-aligned commonsense offers more grounded cues for inference. However, category-based commonsense (CS-PE, CS-EC, and CS-SI) exhibits dataset-dependent effectiveness. On ScienceQA, CS-PE is most beneficial, while on TextVQA, CS-EC dominates, and MMMU shows a more balanced pattern. These results align with our earlier commonsense distribution analysis in Section 4.2, highlighting the importance of tailoring knowledge retrieval to the dataset’s unique characteristics.
@@ 5.2 Implicit Commonsense Knowledge
Table 3: Quantitative analysis on the effect of each component of MAGIC-VQA on the model performance. The "Ex-CS" ($CS$ and $Rel$) denotes explicit commonsense knowledge inclusion (All-CS in Section 5.1), while "Im-CS" ( $Conf$) denote implicit commonsense inclusion. Green check () denotes the inclusion of a component, and red cross () denotes exclusion.
We next examined the effect of implicit commonsense knowledge augmentation as outlined in Section 3.3 using two representative LVLMs, Qwen2VL andGPT4o-mini. We also include results from explicit commonsense knowledge integration (All-CS in Section 5.1). As Table 3 suggests, while the implicit commonsense knowledge $Conf$ does not contribute as significantly as explicit commonsense knowledge (Ex-CS), it nonetheless provides complementary information that enhances overall performance. Incorporating only $Conf$ with Qwen2VL improves the MMMU accuracy from 51.10% to 53.41%, and withGPT4o-mini, the accuracy increases from 55.89% to 57.64%. We also observe that adding $Rel$ notably improves results across all three datasets. Furthermore, combining implicit and explicit commonsense yields the highest overall performance, indicating that implicit augmentation complements explicit knowledge by capturing additional nuances and context that explicit methods alone may miss. Further detailed qualitative analysis is provided in Section 6 and the Appendix B.
@@ 5.3 Break Down Results
We compare performance on specific subcategories within each selected dataset using Qwen2VL and GPT4o-mini in order to analyze the effects of commonsense knowledge augmentation in Figure 4. Across all datasets and subcategories, incorporating commonsense significantly improves the accuracy. Each dataset features distinct subcategories that would benefit from varying aspects of commonsense reasoning.
First, ScienceQA in Figure 4(a), commonsense augmentation yields notable improvements, particularly in language-related subcategories, reflecting the value of context-sensitive reasoning. TextVQA in Figure 4(b), categories involving concrete objects, such as ‘uniform’ and ‘books’, benefit more significantly from commonsense augmentation compared to abstract categories like ‘persons’, indicating that concrete objects allow for more precise retrieval of relevant commonsense knowledge. For MMMU in Figure 4(c), commonsense augmentation benefits easy-level questions, closely tied to everyday knowledge, while struggles with hard-levels that demand complex reasoning beyond commonsense.
5.4 Further Ablation Studies
Beyond our primary experiments, we conducted additional ablation studies to further analyze the impact of various design choices and methodological components. The detailed results for these studies are available in Appendix A. Firstly, as in Appendix A.1, we examined the effect of varying the selection ratio of different commonsense knowledge types—CS-PE, CS-EC, and CS-SI—when integrated with Qwen2VL on the ScienceQA. This analysis helps to determine how the balance of different types of commonsense knowledge influences model performance. Secondly, in Appendix A.2, we investigated the impact of incorporating different numbers of commonsense knowledge triplets using both Qwen2VL and GPT4o-mini. This study aims to assess whether increasing the number of commonsense knowledge triplets enhances reasoning capabilities or if there is a saturation point beyond which performance gains plateau. Thirdly, Appendix A.3 presents the effect of diverse similarity metrics by comparing Manhattan, Cosine, and Euclidean Distance when applied with GPT4o-mini, providing insights into how different similarity measures affect retrieval effectiveness. Finally, as presented in Appendix A.4, we conducted an evaluation of VLPM-style fine-tuning by comparing our proposed approach with methods that distill implicit commonsense knowledge using graph-bassed techniques into compact VLPMs, such as ViLT and ALBEF. This comparison highlights the advantages of our method in effectively integrating commonsense reasoning within smaller VLPMs while maintaining performance efficiency.

### v2

@@ Abstract
Visual Question Answering (VQA) requires reasoning across visual and textual modalities, yet Large Vision-Language Models (LVLMs) often lack integrated commonsense knowledge, limiting their robustness in real-world scenarios. To address this, we introduce MAGIC-VQA, a novel framework that enhances VQA by systematically integrating commonsense knowledge with LVLMs. MAGIC-VQA employs a three-stage process: (1) Explicit Knowledge Integration from external sources, (2) By-Type Post-Processing for contextual refinement, and (3) Implicit Knowledge Augmentation using a Graph Neural Network (GNN) for structured reasoning. While GNNs bring greater depth to structured inference, they enable superior relational inference beyond LVLMs. MAGIC-VQA bridges a key gap by unifying commonsensse knowledge with LVLM-driven reasoning, eliminating the need for extensive pre-training or complex prompt tuning. Our framework achieves state-of-the-art performance on benchmark datasets, significantly improving commonsense reasoning in VQA. Our implementation is open-sourced on GitHub [^1] https://github.com/adlnlp/magic_vqa
@@
Shuo Yang1, Soyeon Caren Han1thanks: *Corresponding author. Caren.Han@unimelb.edu.au, Siwen Luo2, Eduard Hovy1 1The University of Melbourne, 2The University of Western Australia
@@ 1 Introduction
Visual Question Answering (VQA) Antol et al. (2015); Goyal et al. (2017); Yue et al. (2024) is a complex task requiring models to understand the interaction between visual inputs and textual queries. In recent years, Large Vision-Language Models (LVLMs) Dai et al. (2023); Wang et al. (2024b); Liu et al. (2024b); Li et al. (2024a); OpenAI (2024b) have made substantial progress in VQA through extensive pre-training on massive image-text datasets and instruction tuning. These models excel at object-level visual recognition and semantic understanding, capturing attributes such as spatial relationships and contextual details.
Nevertheless, LVLMs often face challenges on questions requiring commonsense reasoning, particularly those hinging on implicit contextual cues or everyday world knowledge Zhou et al. (2023); Ye et al. (2022); Li et al. (2023b) [^2] The sample illustrations can be found in Section 6 and Appendix B.. To overcome this limitation and improve the performance on , different methods have been explored. For example, Multimodal retrieval-augmented generation leverages dense retrieval to inject external multimodal information into the generation process, thereby enhancing the factual grounding of LVLMs Lin and Byrne (2022); Hu et al. (2023). Multimodal prompt tuning harnesses the model’s innate commonsense knowledge by carefully crafting prompts that combine visual and textual cues from representative samples, guiding LVLMs to leverage their internal reasoning for context-rich answers Wei et al. (2022); Zhang et al. (2023). However, static prompt design usually lacks the dynamic adaptability required for novel scenarios, resulting in limited generalization to unseen or diverse inputs. Additionally, graph-based approaches utilize Graph Neural Networks (GNNs) to incorporate structured commonsense knowledge Ravi et al. (2023); Wang et al. (2022), which surpasses the limitations of purely parametric LVLMs, enabling models to capture explicit and implicit knowledge connections via structured graphs.
@@ 2.1 VLPM and LVLMs on VQA
Vision-Language Pretrained Models (VLPMs) like ViLBERT Su et al. (2019), ALBEF Li et al. (2021) and VILT Kim et al. (2021) have advanced Visual Question Answering (VQA) by improving the alignment between visual and textual modalities in the last few yearsLong et al. (2022a). Recently, Large Vision-Language Models (LVLMs) like InstructBLIP Dai et al. (2023), LLaVA Liu et al. (2024b), GPT4o OpenAI (2024b) and Gemini1.5 Gemini Team (2024) further push the boundary of VQA with strong in-context learning capability through extensive pre-training and instruction-tuning on large-scale image-text datasets. While these models have demonstrate their effectiveness in domains such as math Jiang et al. (2024); Wang et al. (2025a), robotics Zhou et al. (2025a), logical reasoning Bi et al. (2025b),medical Wang et al. (2025e, d) and science Lu et al. (2022), these models still face challenges with questions requiring commonsense knowledge that is intuitive and straightforward for humans, such as reasoning based on implicit contextual cues or general world knowledge Ye et al. (2023); Jiang et al. (2023); Chen et al. (2024a); Yang et al. (2025); Zhou et al. (2025b); Bi et al. (2025a); Jiang et al. (2025). The resource-intensive nature of these models further makes it infeasible to train a model from scratch for enhanced commonsense understanding Bi et al. (2024).
@@ 2.2 Commonsense Knowledge Integration for Visual Question Answering
Several studies have highlighted the critical role of commonsense knowledge integration in enhancing the performance of VLPMs and LVLMs on VQA tasks Wu et al. (2022); Zhang et al. (2022); Ding et al. (2022); Wang et al. (2024c); Li et al. (2025a); Wang et al. (2025b). These methods can be classified into two approaches: explicit commonsense knowledge integration and implicit commonsense knowledge integration.
1) Explicit commonsense knowledge integration directly incorporates external commonsense knowledge into model training through instruction tuning, prompt tuning, or reinforcement learning. For example, VLC-BERT Ravi et al. (2023) encodes the contextualized commonsense knowledge of the question phrases as additional textual features and integrates with object visual features to fine-tune the VL-BERT Su et al. (2019). MM-CoT Zhang et al. (2023), T-SciQ Wang et al. (2024a) and KAM-CoT Mondal et al. (2024) fine-tune models on commonsense-augmented Chain-of-Thought (CoT) data to enhance their reasoning processes. Prism Bi et al. (2025a) incorporate the filtered commonsense knowledge into model via dynamic data selection. Vl-Rethinker Wang et al. (2025b) and Pixel Reasoner Su et al. (2025) inject commonsense knowledge into vision-language model via reinforcement learning. However, these methods suffer from static commonsense integration without dynamic filtering to adjust to varying input contexts, resulting in potential noise that impedes nuanced reasoning Li et al. (2024b).
2) Implicit Commonsense Knowledge Integration focuses on transferring knowledge from a teacher to a student model without directly incorporating external datasets. For example, Dai et al. (2022) distill knowledge from the dual-stream CLIP Radford et al. (2021) into BART Lewis (2019), achieving strong zero-shot performance. Park et al. (2024) proposed a novel method to distill knowledge from LLMs on specific image regions, then guiding the LLM to infer commonsense knowledge about those areas. However, these methods often overlook the structured interplay among visual, textual, and commonsense cues, limiting their ability to perform nuanced reasoning Wu et al. (2025).
@@ 3.4 Commonsense Grounded Inference
In the final inference stage, we combine all processed elements—original inputs ($I,Q,C$), refined commonsense triplets (with assigned relevance levels) and GNN-generated confidence scores—into a unified input structure for inference with LVLMs. A complete input example is provided in Appendix F. By fusing explicit and implicit commonsense knowledge with visual and textual signals, the LVLMs can reason effectively about nuanced relationships, delivering answers better aligned with real-world understanding.
@@ 4.1 Dataset
Figure 3: The distribution of categories of commonsense. CS-PE refers to physical entity-related commonsense, CS-EC to event-centered related commonsense, and CS-SI to social interaction-related commonsense.
@@ 4.2 Commonsense Knowledge Distribution
To tailor the commonsense knowledge to each dataset’s specific reasoning requirements, we analyze the distribution of commonsense types across each dataset using GPT4 OpenAI (2024a) [^4] The prompt template is in Appendix G. As Figure 3 suggests, ScienceQA requires more Physical Entity (CS-PE) knowledge, possibly due to its focus on factual scientific concepts. Meanwhile, TextVQA, which often involves contextual understanding in images, benefits more from Event-Centered (CS-EC) knowledge. MMMU, however, requires a balanced mix of Physical Entity, Event-Centered, and Social Interaction (CS-SI) commonsense due to its multi-disciplinary nature. As a result, we set the by-type filtering ratio of {CS-PE:CS-EC:CS-SI} mentioned in Section 3.2 as {0.7:0.15:0.15} in ScienceQA, {0.2:0.6:0.2} in TextVQA,{0.33:0.33:0.33} in MMMU dataset.
@@ 5.1 Explicit Commonsense Knowledge
We evaluated the explicit integration of commonsense knowledge triplets by systematically testing four configurations of: (1) None: Inputs with no additional commonsense; (2) CS Sources: Inputs augmented with commonsense from questions, images, or captions; (3) CS Categories: Inputs augmented with commonsense grouped by category (Physical Entities, Event-Centered, Social Interaction); and (4) All CS: Inputs augmented with all retrieved commonsense [^5] Each experiment is tested with a fixed number of $k=6$ to maintain a fair comparison. We further investigated the impact of incorporating different numbers of commonsense knowledge triplets in Appendix A.1.. As in Table 2, integrating explicit commonsense consistently improves performance across all baselines and three datasets. For instance, on ScienceQA, GPT-4O’s accuracy rises from 76.45% (None) to 81.22% (All CS), and Qwen2VL improves from 51.10% to 57.42% on MMMU under the same setup. Examining the effect of source-based commonsense reveals that image-driven knowledge (CS-I) typically provides the largest gains. For example, LLaVA1.6 on MMMU jumps from 48.38% to 53.52% with CS-I, surpassing the minor improvements from CS-Q or CS-C. This suggests that leveraging image-aligned commonsense offers more grounded cues for inference. However, category-based commonsense (CS-PE, CS-EC, and CS-SI) exhibits dataset-dependent effectiveness. On ScienceQA, CS-PE is most beneficial, while on TextVQA, CS-EC dominates, and MMMU shows a more balanced pattern. These results align with our earlier commonsense distribution analysis in Section 4.2, highlighting the importance of tailoring knowledge retrieval to the dataset’s unique characteristics.
@@ 5.2 Implicit Commonsense Knowledge
Table 3: Quantitative analysis on the effect of each component of MAGIC-VQA. The "Ex-CS" ($CS$ and $Rel$) denotes explicit commonsense knowledge inclusion (All-CS in Section 5.1), while "Im-CS" ( $Conf$) denote implicit commonsense inclusion. () and () denotes the inclusion and exclusion of a component.
We next examined the effect of implicit commonsense knowledge augmentation as outlined in Section 3.3 using two representative LVLMs, Qwen2VL and GPT4o-mini. We also include results from explicit commonsense knowledge integration (All-CS in Section 5.1) as a reference. As Table 3 suggests, while the implicit commonsense knowledge $Conf$ does not contribute as significantly as explicit commonsense knowledge (Ex-CS), it nonetheless provides complementary information that enhances overall performance. Incorporating only $Conf$ with Qwen2VL improves the MMMU accuracy from 51.10% to 53.41%. We also observe that adding $Rel$ notably improves results across all three datasets. Furthermore, combining implicit and explicit commonsense yields the highest overall performance, indicating that implicit augmentation complements explicit knowledge by capturing additional context that explicit methods alone may miss. Further analysis is provided in Section 6 and Appendix B.
Model Knowledge A-OKVQA VCR VLCBERT COMET $38.05$ $79.24$ KAT Wikidata + GPT3 $49.74$ $83.18$ KRISP Wikidata + CNet $27.10$ $65.12$hdashlineOurs (LLaVA1.6) ATOMIC2020 $73.33$ $89.91$ Ours (Qwen2VL) ATOMIC2020 $75.95$ $91.24$ Ours (GPT4o-mini) ATOMIC2020 76.55 93.42
Table 4: Comparison of knowledge-augmented VQA models on two benchmarks. CNet stands for the Conceptnet dataset
5.3 Comparison With Other Baselines
To further assess the effectiveness of MAGIC-VQA’s knowledge injection, we compare it with three representative knowledge-augmented baselines including KAT Gui et al. (2021), VLC-BERT Ravi et al. (2023) and KRISP Marino et al. (2021) on two knowledge-intensive VQA benchmarks: VCR Zellers et al. (2019) and the multiple-choice questions subset of A-OKVQA Schwenk et al. (2022). As shown in Table 4, MAGIC-VQA consistently outperforms all competitors on both datasets. In particular, the GPT-4o-mini variant achieves the highest accuracy of 76.55% on A-OKVQA and 93.42% on VCR, exceeding KAT by more than 26% and 10%, respectively. This improvement highlight the strength of our commonsense-infused visual reasoning. Moreover, unlike prior knowledge-augmented models that require specialized pre-training, MAGIC-VQA is a plug-and-play framework that can be seamlessly integrated with different backbones, making it easily extensible to future models and knowledge bases.
5.4 Ratio of Commonsense Knowledge Type
To determine the optimal ratio of commonsense knowledge types, we conduct an experiment to analyze how varying distributions of CS-PE, CS-EC, and CS-SI knowledge triplets impact the performance of Qwen2-VL on the ScienceQA dataset. Figure 4 indicates that the model achieves its highest accuracy of 75.11% with a 4:1:1 ratio, which places greater emphasis on CS-PE knowledge. Additionally, distributions favoring CS-PE triplets consistently result in improved performance. For example, a 3:1:2 ratio achieves an accuracy of 74.73%. In contrast, ratios prioritizing CS-EC or CS-SI, such as 1:2:3 or 1:3:2, yield lower accuracies of 72.56% and 72.41%, respectively. These results suggest that CS-PE is the most essential commonsense knowledge type for the ScienceQA, aligning with its focus on physical concepts and entities as discussed in Section 5.1. Notably, the 4:1:1 ratio closely mirrors the inherent distribution of commonsense knowledge in ScienceQA in Figure 3. This alignment suggests tailoring the balance of commonsense knowledge to match the dataset’s inherent characteristics, as demonstrated by our approach, leads to the most significant performance improvements. We further attach the effect of diverse similarity metrics by comparing Manhattan, Cosine, and Euclidean Distance when applied to the explicit knowledge retrieval in Appendix A.2.
@@ 5.5 Break Down Results
We provide break down result on subcategories within each selected dataset using Qwen2VL and GPT4o-mini as illustrated in Figure 5. Across all datasets and subcategories, incorporating commonsense significantly improves the accuracy. Each dataset features distinct subcategories that would benefit from varying aspects of commonsense reasoning. First, as illustration of ScienceQA in Figure 5(a), commonsense augmentation yields notable improvements in language-related subcategories, reflecting the value of context-sensitive reasoning. Next shown in Figure 5(b) for TextVQA, categories involving concrete objects such as ‘uniform’ and ‘books’, benefit more from commonsense augmentation compared to abstract categories like ‘persons’, indicating concrete objects allow for more precise retrieval of relevant knowledge. MMMU in Figure 5(c), commonsense augmentation benefits easy-level questions closely tied to everyday knowledge, while struggles with hard-levels that demand complex reasoning beyond commonsense.
5.6 Computation Resource Comparison
We provide a computational resources analysis to demonstrate MAGIC-VQA’s exceptional efficiency and scalability. As highlighted in Table 5, our entire framework spans only 0.33 M trainable parameters, vastly leaner than conventional vision–language transformers, knowledge-augmented transformers, multimodal chain-of-thought models, and today’s large VLMs. This dramatic reduction stems from MAGIC-VQA’s plug-and-play architecture: explicit knowledge retrieval and rule-based filtering are parameter-free, while the implicit commonsense module trains only a lightweight GNN. Unlike previous work that must pre-train or fine-tune hefty backbones to absorb external knowledge, our design cleanly decouples knowledge acquisition from model capacity, enabling rapid adaptation to new vision–language encoders with negligible computational overhead.
Category / Model Year Size Vision–Language Transformers         VL-BERT Su et al. (2019) 2019  [^1] M         ViLT Kim et al. (2021) 2021 112M         UnifiedQA Khashabi et al. (2020) 2020 223M Knowledge Augmented Transformers         KRISP Marino et al. (2021) 2020 116M         KAT Gui et al. (2021) 2021 175B         VICBERT Ravi et al. (2023) 2022 118M Multimodal CoT Models         MM-COT Zhang et al. (2023) 2023 223M         KAM-COT Mondal et al. (2024) 2024 250M Large Vision-Language Models         LLaVA Liu et al. (2023) 2023 7B         GPT4 OpenAI (2024a) 2024 >175B         Qwen2-VL Wang et al. (2024b) 2024 7B Ours         MAGIC-VQA 2024 0.33M
Table 5: Comparison with notable previous methods from different categories with respect to parameter size.